In a survey of 8000 ​women, 5431 say they change their nail polish once a week. Construct a 90% confidence interval for the population proportion of women who change their nail polish once a week.
A 90​% confidence interval for the population proportion is? Round to three decimal places as​ needed

Answer by Boreal(15235)

One sample proportion
90% CI is z* SE
z=1.282
SE=sqrt{p*(1-p))/n)
p=0.6789
1-p=0.3211
n=8000
SE=0.0052
CI on one side=1.282*0.0052=0.0067
90% CI=(0.672,0.686), rounding at the end. This means that we are 90% confident that the true value lies in this interval.